three cans of beans and five cans of soup weigh 700g.


four cans of beans weigh 300g.


what is thr weight of five cans of beans and three cans of soup?.


1
Expert's answer
2022-03-17T06:27:52-0400

Since 4 cans of beans weigh 300 g, each can of beans weigh "\frac{300}{4}=75" g.

Then 5 cans of soup weigh "700-3\cdot 75=475" , from which each can of soup weigh "\frac{475}{5}=95" g.

The weight of 5 cans of beans and 3 cans of soup is "5\cdot 75+3\cdot 95=660" g